In a certain school, the probabilities of the number of students who are reported tardy are shown in the following table:

Number tardy: 0 1 2 3 4
Probability: 0.16 0.25 0.31 0.21 0.07

What is the expected number of tardies (rounded to two decimal places)?
Answer #1

Solution

E(X) = X * P(X)

= 0 * 0.16 + 1 *0.25 + 2 *0.31 + 3 * 0.21 + 4 * 0.7

= ( 0+ 0.25 + 0.62 +0.63+2.8 )

= 4.3